OpenStudy (anonymous):

use type 2 to set up and then evaluate the integration of int_int_ xe^y dA, where D is bounded by y = 2x, y = x^2 anyone help me,please

OpenStudy (anonymous):

@Spacelimbus guide me, please

OpenStudy (anonymous):

Did you draw the regions? Unfortunately I don't have any drawint utilities here at the moment, but basically the region should be all you care about at the moment.

OpenStudy (anonymous):

yes 

OpenStudy (anonymous):

that's the graph I got from the problem

OpenStudy (anonymous):

This might take some fidgeting around with the bounds or rather with the order of integration, but I believe the first we might want to do is set y as constant and integrate with respect to x, that would mean \[\frac{ y }{ 2 } \leq x \leq \sqrt{y}\]

OpenStudy (anonymous):

yes, friend

OpenStudy (anonymous):

and 0<= y<=4 . that's what i got

OpenStudy (anonymous):

very good yes.

OpenStudy (anonymous):

:) but I stuck at take integral. shame on me, cannot get the right answer

OpenStudy (anonymous):

ok I will take a look at in

OpenStudy (amistre64):

x = 0 to 1 y = x^2 to x

OpenStudy (anonymous):

it should turn out to be \[\frac{ 1 }{ 8 }y^2e^y-\frac{ 1 }{ 2 }ye^y\]

OpenStudy (amistre64):

\[\int_{0}^{1}\int_{x^2}^{x}~xe^y~dydx\] or \[\int_{0}^{1}\int_{y}^{\sqrt{y}}~xe^y~dxdy\]

OpenStudy (anonymous):

@amistre64 , the line is y=2x

OpenStudy (amistre64):

... im blinder than a bat :/

OpenStudy (anonymous):

won't change much though!

OpenStudy (anonymous):

dividing by two here, adding 3 there :D

OpenStudy (anonymous):

What are both you talking about? @amistre64 I don't get what you mean @Spacelimbus I got what you got but 1/2 ye^y dy - 1/8 y^2e^ydy

OpenStudy (anonymous):

yes @Hoa you are right, you got that right, I messed up the order when typing it into my equation editor.

OpenStudy (anonymous):

and I stuck at the second part, don't know how to take int. the first part is ok, I got e^4 at the end up. try to solve the leftover.

OpenStudy (anonymous):

well these are two integrals you can solve with the tabular method for integration, have you ever done that @Hoa

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

teach me, please. it sounds new to me

OpenStudy (anonymous):

Lets check the more complicated one.